the position vector of a particle is given by r=2t i^-5t^2 j^.what is the angle between initial velocity and initial acceleration

Grade:11

1 Answers

Vidhi Shah
6 years ago
x=2t=>vx=dxdt=2y=2t2=>vy=dydt=4ttanθ=vyvx=4t2=2tDifferentiating with respect to t(sec2θ)dθdt=2(1+tan2θ)dθdt=2(1+4t2)dθdt=2dθdt=21+4t2at t=2dθdt=21+4(2)2=217rad/sHence a is the correct answer.